5. Twenty-nine randomly selected women had their Body Mass Indices (BMI) measured. The sample standard deviation was 7.6 poinPlease Show work. Thanks.
Answer #1

Date Page s=7.6 solution Given that, n=29 (1-4)% 80 -0.80 x = -0.80 = 0.20 2/2 = 0.10 1-4/2= 10.10 - 0.90 The critical valuesDate Page (n-1) 52 (-)) 22 < < 22 X2 - ر4 / 1 (8) راد (۱) 2 (26) (۱-مد) (ی.) (۱- ۹د) که 4 37.416 18. 939 28 (76) 2 31.94 (( 26.5 < 5 29.2 The 80% confidence interval for the population standard deviation lies between the C6.59.2)